# Can AR Embedded Visualizations Foster Appropriate Reliance on AI in Spatial Decision-Making? A Comparative Study of AR X-Ray vs. 2D Minimap

### v2

@@ Abstract.
Artificial Intelligence (AI) and indoor sensing increasingly support decision-making in spatial environments. However, traditional visualization methods impose a substantial mental workload when viewers translate this digital information into real-world spaces, leading to inappropriate reliance on AI. Embedded visualizations in Augmented Reality (AR), by integrating information into physical environments, may reduce this workload and foster more appropriate reliance on AI. To assess this, we conducted an empirical study (N = 32) comparing an AR embedded visualization (X-ray) and 2D Minimap in AI-assisted, time-critical spatial target selection tasks. Surprisingly, evidence shows that the embedded visualization led to greater inappropriate reliance on AI, primarily as over-reliance, due to factors like perceptual challenges, visual proximity illusions, and highly realistic visual representations. Nonetheless, the embedded visualization demonstrated benefits in spatial mapping. We conclude by discussing empirical insights, design implications, and directions for future research on human-AI collaborative decision in AR.
@@
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3790710isbn: 979-8-4007-2278-3/2026/04
Figure 1. A participant wearing an Apple Vision Pro performs spatial decision-making tasks using two different visualizations: AR X-ray (Left) and 2D Minimap (Right). Both visualizations present AI-suggested targets, but the AR X-ray approach embeds information directly into the participant’s field of view, while the 2D Minimap offers a top-down abstraction. The depiction of targets across two floors is for illustration only and does not reflect the actual experimental setup.
@@ 1. Introduction
Augmented Reality (AR) promises a more intuitive approach by directly visualizing data and AI suggestions onto the physical environment. These visualizations, often referred to as embedded visualizations (Willett2017EmbeddedDataRepresentations; lee2023design; Bressa2022SituatedVisualization; Shin2024TheRealityofTheSituation), can reduce cognitive load by eliminating the need for constant reference frame translation. Intuitively, embedded visualizations might foster more deliberate and appropriate reliance on AI. Yet, whether embedded visualizations genuinely mitigates inappropriate reliance on AI in spatial decision-making remains unclear, and little existing work has examined how embedded visualizations in AR affects human-AI reliance in spatial tasks. As AR and AI continue to mature and increasingly integrate (Hirzle2023WhenXRAndAIMeet), this work aims to provide researchers and practitioners with initial evidence on how embedded visualizations impact human-AI reliance in spatial decision-making.
@@
To this end, we present an empirical study (N=32) comparing an AR X-ray (Figure 1 Left) against a 2D Minimap (Figure 1 Right) in a time-pressured, AI-assisted spatial decision task. Following previous work on AR visual systems for indoor navigation (XU2021Seethroughwalls; XUImroveIndoorWayfinding; XUARteambasedsearch), we selected the X-ray view as our AR embedded visualization because it naturally embeds information within a large-scale indoor environment. In the tasks, participants needed to select one of four targets distributed across a two-floor building based on multiple criteria under time pressure. An AI with 75% accuracy was provided to suggest the optimal target, and both visualizations were experienced using an Apple Vision Pro (AVP)  [^1] https://www.apple.com/apple-vision-pro/, a state-of-the-art AR headset. This abstract task models time-pressured spatial choices with imperfect AI support, where over-reliance can hide errors and under-reliance can waste useful guidance.
@@
Contrary to our expectations, participants exhibited greater inappropriate reliance on AI in the X-ray condition, suggesting that new perceptual and cognitive challenges arise when digital cues are spatially embedded in large-scale physical environments. Further analysis of quantitative data (decision accuracy, response times, reliance metrics, and self-reported confidence) reveals that inappropriate reliance manifested primarily as over-reliance (i.e., blindly accepting AI suggestions). Through qualitative interviews, we identify several key factors that contributed to the over-reliance, such as perceptual challenges in AR, visual proximity illusions, and heightened trust in “embodied” AI suggestions. Nonetheless, our findings suggest the X-ray’s strength in spatial mapping (as evidenced by quantitative data) and support for egocentric spatial imagery (as reported by participants). These findings matter for AR+AI design because they show that while embedding AI cues into the physical world can strengthen spatial mapping, it can also simultaneously miscalibrate reliance. We conclude by outlining the lessons learned, design implications, and promising future directions for better harnessing these AR strengths to foster human–AI collaboration in spatial decision-making.
@@ 2.2. Visualization-Aided Decision-Making
Within the field of visualization, research on decision-making typically falls into two major categories: understanding how humans use visualizations to make decisions (padilla2018decision; dimara2018task; bearfield2024same) and design interactive systems (Oral_2023_ACriticalInquiry).
@@ 2.3. Spatial Cognition and Spatial Decision-Making
Spatial cognition encompasses the mental processes involved in perceiving, encoding, and reasoning about spatial environments (Tversky_2008_SpatialCognition). A fundamental aspect of spatial cognition is the use of reference frames—mental models that help individuals locate and orient themselves in space (Tversky_2008_SpatialCognition; Barbara2003StructureOfMentalSpace). Two commonly described reference frames are egocentric, where objects are located relative to the observer’s current position and orientation (i.e., “to my left or right”), and allocentric, where the environment is represented independently of the observer’s position, such as map-like or object-to-object relationships (Klatzky1998AllocentricAE). People switch flexibly between these frames, but doing so can introduce substantial cognitive load (Barbara2003StructureOfMentalSpace), particularly when tasks require precise mental transformations, such as rotating or scaling spatial representations in mind.
@@ 2.4. AR, Situated, and Embedded Visualizations
Different from Virtual Reality (VR) (burdea2003virtual), AR can directly visualize information within physical environments, benefiting a broad range of real-world applications that require computational support (Kalkofen2009ComARVis). The visualization community has explored AR visualizations across multiple dimensions, including design patterns (Shin2024TheRealityofTheSituation; lee2023design; Bressa2022SituatedVisualization), techniques (patnaik2024vistorch; Zhu_Tian_2024), systems (fleck2022ragrug; liu2025reality), empirical evaluations (quijano2024brushing; de2019defamiliarization), and applications (Lin_2021; lee2024sportify). We refer readers to these recent surveys (Bressa2022SituatedVisualization; Shin2024TheRealityofTheSituation; lee2023design) for more details.
More broadly, the concept of visualizing information in its physical context is known as Situated Visualization (Shin2024TheRealityofTheSituation; lee2023design; Bressa2022SituatedVisualization), for which AR is one prominent realization among others like physicalization (Jansen2015DataPhysical). Depending on the spatial relationship between the visualization and its physical referent, Willett et al. (Willett2017EmbeddedDataRepresentations) distinguish between situated (co‑located in physical space) and embedded (overlaid or integrated with the referent). Based on this definition, our study compares two AR visualizations: a situated 2D Minimap and an embedded AR X-ray view. This comparison enables us to investigate whether embedding data directly into the physical referent can foster more appropriate reliance on AI by reducing the mental effort required to map data from a separate panel onto the physical environment. For clarity in the remainder of the paper, we refer to these two AR visualizations simply as Minimap and X-ray, respectively.
@@ 3.1. Task Abstraction and Definition
Spatial decision-making has long been studied under the broader umbrella of wayfinding in spatial cognition (Emo2012WayfindingAS; Wiener2009TaxonomyOH). Foundational work by Siegel and White (Siegel1975TheDO) conceptualizes wayfinding as a decision process involving three forms of spatial knowledge: landmark, route, and survey. Motivated by real-world applications, our study particularly focuses on target selection (i.e., landmark). In tasks such as exploring a building, security surveillance, and emergency response, spatial decision-making often involves identifying or selecting a target before proceeding with an action like navigation or intervention. Although AR systems have frequently been investigated for supporting the action phase such as route guidance (XUARteambasedsearch), their potential to enhance the preceding decision-making step remains underexplored.
@@
  • Data - Static and Dynamic: Appropriate target selection relies on both static data like building level layouts and dynamic updates (e.g., walking people, occupant counts, hazard alerts) from sensors to evaluate each target (Adib2013SeeThroughWallsWithWifi).
@@
X-Ray. Given our task focuses on large-scale indoor environments, we select AR X-ray (Figure 1 Left) as our embedded visualizations, which overlays a 1:1 scale 3D map on the real-world. It allows users see the interior layouts or objects behind walls, as if they had “X-ray” vision (Livingston2013PursuitO), which has been widely used in applications like navigation (XUImroveIndoorWayfinding) and construction (Soria2018AugmentedAV). For our study, we follow previous similar research (XU2021Seethroughwalls; XUARteambasedsearch; XUImroveIndoorWayfinding) and portray the building’s indoor layout, along with multiple targets (represented as cubes), and their associated data (i.e., the number of people in line) using virtual avatars.
@@ 4.3. Implementation and Apparatus
Following similar work (2021ToTrustOrToThink), a simulated AI was used to provide decision suggestions in applicable trials. Details of the simulated AI setup are provided in Sec. 4.7. We developed the minimap based on high-precision CAD models provided by the building manager. The AR visualizations was developed using ARKit  [^3] https://developer.apple.com/augmented-reality/arkit/ and AR Mesh Manager  [^4] https://docs.unity3d.com/Packages/com.unity.xr.arkit@6.2/manual/arkit-meshing.html to scan the building and construct a 3D digital twin in Unity 6000.0.29f1, with ARAnchor Manager  [^5] https://docs.unity3d.com/Packages/com.unity.xr.arfoundation@6.2/manual/features/anchors/aranchormanager.html ensuring proper registration of the digital model with the physical environment. The system is available at our public repository http://public-when-ready. All trials were completed using an Apple Vision Pro to eliminate device-related differences and provided a high-fidelity immersive experience.
@@ 32 Participants
  • Actual Tasks (30mins): Participants completed 32 trials, grouped into four blocks, with each block distributed over 2 consecutive locations (see the example in Figure 5). Within each block, one condition was tested (i.e., eight trials per condition). Within each location, they remained physically stationary and completed four trials (one trial per difficulty level). Breaks were allowed during the study.
@@ 5.1.2. AI Reliance
To interpret these numbers, we compared the observed behaviors with a random-selection baseline. Among the 8 trials in each block, the AI suggested 5 optimal, 2 sub-optimal, and 1 worst target. If a participant were to choose a target at random (i.e., with equal probability of $1/4$ for each of the four possible targets), the expected numbers of over- and under-reliance trails per block are 0.75 and 4.25, respectively (see Supplemental Materials for calculation details). Compared to this baseline, participants in the X-ray+AI condition displayed fewer instances of under-reliance ($M_{\text{{X-ray}}}=1.63<M_{\text{random}}=4.25$), but notably more instances of over-reliance ($M_{\text{{X-ray}}}=1.41>M_{\text{random}}=0.75$). In contrast, participants in the Minimap+AI condition exhibited fewer instances of both under-reliance ($M_{\text{{Minimap}}}=0.84<M_{\text{random}}=4.25$) and over-reliance ($M_{\text{{Minimap}}}=0.72<M_{\text{random}}=0.75$). These findings indicate that inappropriate reliance in the X-ray+AI condition was primarily driven by over-reliance, consistent with our earlier accuracy analysis (Sec. 5.1.1).
@@ 5.1.3. Response Time
Response Time Excluding Initial Search. We observed that participants in the X-ray conditions spent additional time to visually search the spatially distributed targets—especially those outside their immediate field of view. In contrast, in the Minimap conditions, all targets were presented in a 2D overview, making search overhead negligible. Prior decision-making models (Simon’s model (Simon1977), the drift–diffusion model (ratcliff2008diffusion), the leaky competing accumulator model (usher2001time)) and empirical studies (reisen2008identifying; russo1994eye) suggest that actual decision-making begins only after sufficient information has been gathered. Thus, similar to prior research (Jakobsen2013VISonDisplay), we distinguished the initial search phase from the subsequent decision-making phase. This motivates us to exclude the initial search overhead in the X-ray conditions to better compare its decision-making phase with the Minimap conditions.
@@
Following prior visualization-decision work (Padilla2018DecisionMW), we consider that meaningful decision-making start only after at least two options have been identified. Thus, for each trial in the X-ray conditions, we excluded the initial search time required to discover the first two distinct targets. Specifically, we defined two search segments to exclude: S1-from the start of the trial to the first frame when the first target entered the user’s field of view; and S2-from the frame when the user turned away from the first target (Figure 9a) to the frame when the second target appeared (Figure 9b). If a target was already visible at trial start, or if the first and second targets appeared simultaneously, the corresponding segment (S1 or S2) was set to zero. This definition provides a strict lower bound on search time, as participants may still require additional head movement to center the target in view. Yet, we acknowledge that this measure is not a theoretically perfect indicator of true decision time. Nevertheless, it provides a scientifically reasonable approximation that is consistent with definitions and procedures used in prior research (Jakobsen2013VISonDisplay; Padilla2018DecisionMW; reisen2008identifying; russo1994eye). We discuss its limitations and possible refinements in Sec. 6.5.

### v3

@@ Abstract.
Artificial Intelligence (AI) and indoor sensing increasingly support decision-making in spatial environments. However, traditional visualization methods impose a substantial mental workload when viewers translate this digital information into real-world spaces, leading to inappropriate reliance on AI. Embedded visualizations in Augmented Reality (AR), by integrating information into physical environments, may reduce this workload and foster more appropriate reliance on AI. To assess this, we conducted an empirical study (N = 32) comparing an AR embedded visualization (X-ray) and 2D Minimap in AI-assisted, time-critical spatial target selection tasks. Surprisingly, evidence shows that the embedded visualization led to greater inappropriate reliance on AI, primarily as over-reliance, due to factors like perceptual challenges, visual proximity illusions, and highly realistic visual representations. Nonetheless, the embedded visualization showed benefits in spatial mapping. We conclude by discussing empirical insights, design implications, and directions for future research on human-AI collaborative decision in AR.
@@
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3790710isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Empirical studies in visualizationccs: Human-centered computing Empirical studies in HCIccs: Computing methodologies Machine learning
Figure 1. A participant wearing an Apple Vision Pro performs spatial decision-making tasks using two different visualizations: AR X-ray (Left) and 2D Minimap (Right). Both visualizations present AI-suggested targets, but the AR X-ray approach embeds information directly into the participant’s field of view, while the 2D Minimap offers a top-down abstraction.
@@ 1. Introduction
Augmented Reality (AR) promises a more intuitive approach by directly visualizing data and AI suggestions onto the physical environment. These visualizations, often referred to as embedded visualizations (Willett2017EmbeddedDataRepresentations), can reduce cognitive load by eliminating the need for constant reference frame translation. Intuitively, embedded visualizations might foster more deliberate and appropriate reliance on AI. Yet, whether embedded visualizations genuinely mitigates inappropriate reliance on AI in spatial decision-making remains unclear, and little existing work has examined how embedded visualizations in AR affects human-AI reliance in spatial tasks. As AR and AI continue to mature and increasingly integrate (Hirzle2023WhenXRAndAIMeet), this work aims to provide researchers and practitioners with initial evidence on how embedded visualizations impact human-AI reliance in spatial decision-making.
@@
To this end, we present an empirical study (N=32) comparing an AR X-ray (Figure 1 Left) against a 2D Minimap (Figure 1 Right) in a time-pressured, AI-assisted spatial decision task. Following previous work on AR visual systems for indoor navigation (XU2021Seethroughwalls), we selected the X-ray view as our AR embedded visualization because it naturally embeds information within a large-scale indoor environment. In the tasks, participants needed to select one of four targets distributed across a two-floor building based on multiple criteria under time pressure. An AI with 75% accuracy was provided to suggest the optimal target, and both visualizations were experienced using an Apple Vision Pro (AVP)  [^1] https://www.apple.com/apple-vision-pro/, a state-of-the-art AR headset. This abstract task models time-pressured spatial choices with imperfect AI support, where over-reliance can hide errors and under-reliance can waste useful guidance.
@@
Contrary to our expectations, participants exhibited greater inappropriate reliance on AI in the X-ray condition, suggesting that new perceptual and cognitive challenges arise when digital cues are spatially embedded in large-scale physical environments. Further analysis of quantitative data (decision accuracy, response times, reliance metrics, and self-reported confidence) reveals that inappropriate reliance manifested primarily as over-reliance, which means, blindly accepting AI suggestions. Through qualitative interviews, we identify several key factors that contributed to the over-reliance, such as perceptual challenges in AR, visual proximity illusions, and heightened trust in “embodied” AI suggestions. Nonetheless, our findings suggest the X-ray’s strength in spatial mapping (as evidenced by quantitative data) and support for egocentric spatial imagery (as reported by participants). These findings matter for AR+AI design because they show that while embedding AI cues into the physical world can strengthen spatial mapping, it can also simultaneously miscalibrate reliance. We conclude by outlining the lessons learned, design implications, and promising future directions for better harnessing these AR strengths to foster human–AI collaboration in spatial decision-making.
@@ 2.2. Visualization-Aided Decision-Making
Within the field of visualization, research on decision-making typically falls into two major categories: understanding how humans use visualizations to make decisions (padilla2018decision) and design interactive systems (Oral_2023_ACriticalInquiry).
@@ 2.3. Spatial Cognition and Spatial Decision-Making
Spatial cognition encompasses the mental processes involved in perceiving, encoding, and reasoning about spatial environments (Tversky_2008_SpatialCognition). A fundamental aspect of spatial cognition is the use of reference frames—mental models that help individuals locate and orient themselves in space (Barbara2003StructureOfMentalSpace). Two commonly described reference frames are egocentric, where objects are located relative to the observer’s current position and orientation like “to my left or right”, and allocentric, where the environment is represented independently of the observer’s position, such as map-like or object-to-object relationships (Klatzky1998AllocentricAE). People switch flexibly between these frames, but doing so can introduce substantial cognitive load (Barbara2003StructureOfMentalSpace), particularly when tasks require precise mental transformations, such as rotating or scaling spatial representations in mind.
@@ 2.4. AR, Situated, and Embedded Visualizations
Different from Virtual Reality (VR) (burdea2003virtual), AR can directly visualize information within physical environments, benefiting a broad range of real-world applications that require computational support (Kalkofen2009ComARVis). The visualization community has explored AR visualizations across multiple dimensions, including design patterns (lee2023design), techniques (Zhu_Tian_2024), systems (liu2025reality), empirical evaluations (quijano2024brushing), and applications (Lin_2021; lee2024sportify). We refer readers to the recent survey (Shin2024TheRealityofTheSituation) for more details.
More broadly, the concept of visualizing information in its physical context is known as Situated Visualization (Bressa2022SituatedVisualization), for which AR is one prominent realization among others like physicalization (Jansen2015DataPhysical). Depending on the spatial relationship between the visualization and its physical referent, Willett et al. (Willett2017EmbeddedDataRepresentations) distinguish between situated (co‑located in physical space) and embedded (overlaid or integrated with the referent). Based on this definition, our study compares two AR visualizations: a situated 2D Minimap and an embedded AR X-ray view. This comparison enables us to investigate whether embedding data directly into the physical referent can foster more appropriate reliance on AI by reducing the mental effort required to map data from a separate panel onto the physical environment. For clarity in the remainder of the paper, we refer to these two AR visualizations simply as Minimap and X-ray, respectively.
@@ 3.1. Task Abstraction and Definition
Spatial decision-making has long been studied under the broader umbrella of wayfinding in spatial cognition (Emo2012WayfindingAS; Wiener2009TaxonomyOH). Foundational work by Siegel and White (Siegel1975TheDO) conceptualizes wayfinding as a decision process involving three forms of spatial knowledge: landmark, route, and survey. Motivated by real-world applications, our study particularly focuses on target selection (landmark). In tasks such as exploring a building, security surveillance, and emergency response, spatial decision-making often involves identifying or selecting a target before proceeding with an action like navigation or intervention. Although AR systems have frequently been investigated for supporting the action phase such as route guidance (XUARteambasedsearch), their potential to enhance the preceding decision-making step remains underexplored.
@@
  • Data - Static and Dynamic: Appropriate target selection relies on both static data like building level layouts, and dynamic updates from sensors like walking people, occupant counts, and hazard alerts to evaluate each target (Adib2013SeeThroughWallsWithWifi).
@@
X-Ray. Given our task focuses on large-scale indoor environments, we select AR X-ray (Figure 1 Left) as our embedded visualizations, which overlays a 1:1 scale 3D map on the real-world. It allows users see the interior layouts or objects behind walls, as if they had “X-ray” vision (Livingston2013PursuitO), which has been widely used in applications like navigation (XUImroveIndoorWayfinding) and construction (Soria2018AugmentedAV). For our study, we follow previous similar research (XUImroveIndoorWayfinding) and portray the building’s indoor layout, along with multiple targets (represented as cubes), and their associated data (the number of people in line) using virtual avatars.
@@ 4.3. Implementation and Apparatus
Following similar work (2021ToTrustOrToThink), a simulated AI was used to provide decision suggestions in applicable trials. Details of the simulated AI setup are provided in Sec. 4.7. We developed the minimap based on high-precision CAD models provided by the building manager. The AR visualizations was developed using ARKit  [^3] https://developer.apple.com/augmented-reality/arkit/ and AR Mesh Manager  [^4] https://docs.unity3d.com/Packages/com.unity.xr.arkit@6.2/manual/arkit-meshing.html to scan the building and construct a 3D digital twin in Unity 6000.0.29f1, with ARAnchor Manager  [^5] https://docs.unity3d.com/Packages/com.unity.xr.arfoundation@6.2/manual/features/anchors/aranchormanager.html ensuring proper registration of the digital model with the physical environment. The system is available at our public repository https://github.com/demoPlz/ARAIReliance. All trials were completed using an Apple Vision Pro to eliminate device-related differences and provided a high-fidelity immersive experience.
@@ 32 Participants
  • Actual Tasks (30mins): Participants completed 32 trials, grouped into four blocks, with each block distributed over 2 consecutive locations (see the example in Figure 5). Within each block, one condition was tested, namely, eight trials per condition. Within each location, they remained physically stationary and completed four trials (one trial per difficulty level). Breaks were allowed during the study.
@@ 5.1.2. AI Reliance
To interpret these numbers, we compared the observed behaviors with a random-selection baseline. Among the 8 trials in each block, the AI suggested 5 optimal, 2 sub-optimal, and 1 worst target. If a participant were to choose a target at random, namely with equal probability of $1/4$ for each of the four possible targets, the expected numbers of over- and under-reliance trails per block are 0.75 and 4.25 respectively (see Supplemental Materials for calculation details). Compared to this baseline, participants in the X-ray+AI condition displayed fewer instances of under-reliance ($M_{\text{{X-ray}}}=1.63<M_{\text{random}}=4.25$), but notably more instances of over-reliance ($M_{\text{{X-ray}}}=1.41>M_{\text{random}}=0.75$). In contrast, participants in the Minimap+AI condition exhibited fewer instances of both under-reliance ($M_{\text{{Minimap}}}=0.84<M_{\text{random}}=4.25$) and over-reliance ($M_{\text{{Minimap}}}=0.72<M_{\text{random}}=0.75$). These findings indicate that inappropriate reliance in the X-ray+AI condition was primarily driven by over-reliance, consistent with our earlier accuracy analysis (Sec. 5.1.1).
@@ 5.1.3. Response Time
Response Time Excluding Initial Search. We observed that participants in the X-ray conditions spent additional time to visually search the spatially distributed targets—especially those outside their immediate field of view. In contrast, in the Minimap conditions, all targets were presented in a 2D overview, making search overhead negligible. Prior decision-making model (Simon’s model (Simon1977)) and empirical studies (reisen2008identifying; russo1994eye) suggest that actual decision-making begins only after sufficient information has been gathered. Thus, similar to prior research (Jakobsen2013VISonDisplay), we distinguished the initial search phase from the subsequent decision-making phase. This motivates us to exclude the initial search overhead in the X-ray conditions to better compare its decision-making phase with the Minimap conditions.
@@
Following prior visualization-decision work (Padilla2018DecisionMW), we consider that meaningful decision-making start only after at least two options have been identified. Thus, for each trial in the X-ray conditions, we excluded the initial search time required to discover the first two distinct targets. Specifically, we defined two search segments to exclude: S1-from the start of the trial to the first frame when the first target entered the user’s field of view; and S2-from the frame when the user turned away from the first target (Figure 9a) to the frame when the second target appeared (Figure 9b). If a target was already visible at trial start, or if the first and second targets appeared simultaneously, the corresponding segment (S1 or S2) was set to zero. This definition provides a strict lower bound on search time, as participants may still require additional head movement to center the target in view. Yet, we acknowledge that this measure is not a theoretically perfect indicator of true decision time. Nevertheless, it provides a scientifically reasonable approximation that is consistent with definitions and procedures used in prior research (Jakobsen2013VISonDisplay). We discuss its limitations and possible refinements in Sec. 6.5.